# BooookScore: A systematic exploration of book-length summarization in the era of LLMs

### v2

@@ Abstract
Summarizing book-length documents (>>100K tokens) that exceed the context window size of large language models (LLMs) requires first breaking the input document into smaller chunks and then prompting an LLM to merge, update, and compress chunk-level summaries. Despite the complexity and importance of this task, it has yet to be meaningfully studied due to the challenges of evaluation: existing book-length summarization datasets (e.g., BookSum) are in the pretraining data of most public LLMs, and existing evaluation methods struggle to capture errors made by modern LLM summarizers. In this paper, we present the first study of the coherence of LLM-based book-length summarizers implemented via two prompting workflows: (1) hierarchically merging chunk-level summaries, and (2) incrementally updating a running summary. We obtain 1193 fine-grained human annotations on GPT-4 generated summaries of 100 recently-published books and identify eight common types of coherence errors made by LLMs. Because human evaluation is expensive and time-consuming, we develop an automatic metric, BooookScore, that measures the proportion of sentences in a summary that do not contain any of the identified error types. BooookScore has high agreement with human annotations and allows us to systematically evaluate the impact of many other critical parameters (e.g., chunk size, base LLM) while saving \$15K and 500 hours in human evaluation costs. We find that closed-source LLMs such as GPT-4 and Claude 2 produce summaries with higher BooookScore than the oft-repetitive ones generated by LLaMA 2. Incremental updating yields lower BooookScore but higher level of detail than hierarchical merging, a trade-off sometimes preferred by human annotators. We release our code and annotations to spur more principled research on book-length summarization [^1] https://github.com/lilakk/BooookScore.
@@ 1 Introduction
Despite the promise that LLMs hold for long-context tasks, the research community still lacks a principled and systematic approach to evaluate their capabilities on book-length summarization. Our paper identifies three open challenges with evaluation: (1) data contamination, in which existing benchmarks such as BookSum (Kryscinski et al., 2022) are in the pretraining data of modern LLMs; (2) an unexplored error distribution, as most prior summarization research centers around short source documents and fails to capture coherence errors that are exacerbated by the “chunk and combine” book-length summarization setting; and (3) a lack of any reliable automatic metric, which requires careful design and validation against human annotations.
@@
To mitigate the impact of data contamination, we design our evaluation framework around the use of newly-published books. We eschew the collection of gold reference summaries (a practice that is neither scalable nor tractable for book-length documents) and instead propose a protocol that leverages human annotation of the coherence of LLM-generated summaries (i.e., their logical connectedness) under different prompting strategies. Our protocol unifies and extends best-practices across disparate works in document understanding and evaluation research, including adoption of fine-grained annotation units (Krishna et al., 2023), use of QA pairs to denote points of confusion (Ko et al., 2020), and a taxonomic breakdown of different coherence errors (Goyal et al., 2022a).
We validate our protocol by collecting 1193 span-level human annotations on GPT-4 generated summaries of a carefully curated set of 100 recently-published books (costing \$3K and 100 annotator hours) using two prompting strategies (hierarchical merging and incremental updating, shown in Figure 1). In categorizing these annotations into eight frequent error types, we reveal an error distribution in GPT-4 summaries that differs from that observed in prior studies on short-document summarizers (Goyal et al., 2022a); notably, we identify new error types (causal omissions, salience errors) through our book-length summarization setting (Table 3).
@@
We use BooookScore to evaluate the impact of several critical design decisions on the coherence of generated summaries, including the choice of prompting strategy, base LLM, and chunk size, a study that altogether cost \$9K in LLM API calls. Our findings include (1) hierarchical merging generally results in more coherent summaries but reduced level of detail compared to incremental updating; (2) GPT-4 and Claude 2 produce the most coherent summaries, while LLaMA 2 is substantially worse and fails to follow instructions; (3) increasing the chunk size does not improve hierarchical merging but does substantially benefit Claude 2 when using incremental updating; and (4) summary-level preference judgments are highly subjective and do not correlate with BooookScore.
All prompts used in this paper can be found in Appendix J.
@@ 2 Background: summarizing book-length texts with LLMs
One major issue with the hierarchical merging approach is that it necessitates summarizing portions of the input document without complete context. For example, in the first level, chunks towards the end of the book will be summarized without knowledge of what came before, which can lead to incoherent summaries especially for non-linear or multi-perspective narratives. We thus explore an alternate prompting strategy—incremental updating (Figure 1, right)— that iterates through each chunk in order while continuously updating a global summary with salient information. While this method may be better able to handle inter-chunk dependencies than hierarchical merging, it requires more complicated prompts for (1) summarizing an input chunk, (2) updating the global summary $s_{1,2,\dots,i-1}$ with information from the current chunk $c_{i}$, and (3) compressing the global summary when it exceeds the maximum summary length $G_{n}$. See Appendix A.2 for a full specification of incremental updating.
@@ 3 Evaluating coherence of book summaries
The only existing public dataset for book-length summarization is BookSum (Kryscinski et al., 2022), which contains famous books from the Project Gutenberg public-domain repository along with reference summaries scraped from popular websites such as CliffNotes and GradeSaver. Both the source books and reference summaries are in the pretraining data of existing LLMs: Chang et al. (2023) confirm that many books in the BookSum held-out split (e.g., The Adventures of Huckleberry Finn, The Picture of Dorian Gray) are among the most-memorized books by GPT-4 and ChatGPT, and we were able to auto-complete several reference BookSum summaries by prompting GPT-4 with a short prefix of the summary.
To reduce the confounding impact of summary memorization, we manually collect 100 books [^4] We roughly balance our dataset across the following genres: fiction, non-fiction, sci-fi, fantasy, historical, contemporary, and memoir. We also include both linear and non-linear (multi-perspective and time-shifting) narratives in the dataset, and we purchase electronic copies of each of the 100 books in the dataset. published within the past year to form our dataset (see Table 3 for a full list). Some of these books could still have appeared in the pretraining dataset of recent LLMs such as Claude 2 and LLaMa2, although it is much less likely than in BookSum. However, summaries of these books do not publicly exist: we did not find summaries online for any books in our dataset, which means LLM memorization of these summaries is impossible. [^5] However, we did find book reviews, which intentionally do not reveal major plot points or other spoilers. The average length of the books in our dataset is 190K tokens, compared to 112K tokens in BookSum. Due to copyright laws, we cannot publicly release this dataset; even if we could, we would still recommend that researchers collect their own datasets of newly-published books to minimize contamination with LLMs of the future.
@@
Since we lack gold summaries, we design our evaluation framework to be reference-free, which aids in scalability. To do this, our evaluation framework synthesizes best-practices of prior document understanding and summarization evaluation research. Our evaluation employs: (1) fine-grained evaluation units as recommended by LongEval (Krishna et al., 2023); (2) information-seeking questions to represent naturally-occurring points of confusion (Ko et al., 2020); and (3) focus on summary coherence, which evaluates the logical structure and readability of the summary itself (Goyal et al., 2022a). We do not directly evaluate the faithfulness of the summaries (i.e., how factually accurate they are at conveying information from the source text), as the length of the source texts poses considerable issues for any existing faithfulness evaluation. We qualitatively discuss faithfulness in Section 5 and leave further investigation for future work.
@@
We implement our framework through a source- and reference-free annotation protocol where (1) annotators read through an LLM-generated summary, (2) highlight all confusing spans, and (3) ask question(s) for each marked span that highlight their confusion. [^6] We also enabled forming relations between two spans in case multiple spans contributed to the same issue. See Table 3 (third column) for examples of spans and questions produced by our annotators. We hired four annotators with extensive English proofreading experience on Upwork [^7] http://upwork.com, each of whom annotated 25 disjoint summaries. Each summary takes roughly 30 minutes to fully annotate with spans and questions, and we paid \$15 per summary for a total of \$3K to evaluate both prompting strategies. To generate the summaries, we set the base LLM to GPT-4 with a chunk size of 4096 and a maximum summary length $G_{n}=1200$; other hyperparameters are detailed in Section 5. In total, the annotators mark 840 coherence errors for GPT-4-generated summaries with incremental updating, compared to 353 errors for hierarchical merging; see Table 3 (right) for the split across error types.
@@
Typical measures of agreement are difficult to obtain in our setup, as it is expensive to obtain multiple overlapping annotations due to the complexity of the task; additionally, Goyal et al. (2022a) and  Dou et al. (2022) observed low recall among annotators when evaluating machine-generated text at a fine-grained level. This motivates us to instead measure the precision of a given error annotation (i.e., after reading the corresponding question, do you agree that the span is confusing?), as it is a simpler and cheaper task. Given a span from a summary marked as containing an error, along with questions highlighting the confusion, we ask annotators (1) whether they think the span is confusing; and (2) whether the corresponding questions highlight the central confusion. We use the same four annotators hired before for this task, but make them validate human and (and later GPT-4) annotations for 25 books that they did not annotate in the first task. Overall, we validated 1,659 annotations for a total cost of \$418.90, [^8] This cost includes validation of both human and BooookScore annotations. and we discover that 79.7% of annotated spans are validated as legitimate through this task. More details on our validation can be found in Appendix G.
@@
Table 1: Definition of all coherence error types and an example annotation for each. The last column shows the % of errors of specific types per sentence for incremental and hierarchical summaries.
@@ 4 BooookScore: an automatic evaluation metric
Since human evaluation of summary coherence is not scalable due to the high financial and time cost, we develop an automatic metric — BooookScore— that prompts an LLM to identify instances of the eight error types we identified in Section 3. We validate BooookScore via a human evaluation of its precision (following the annotation task discussed in the previous section) and show that its precision matches that of human annotators (78.2% vs. 79.7%). We then use BooookScore to evaluate many other book-length summarization configurations, saving \$15K in evaluation costs and 500 hours in annotator time. We emphasize that incorporating definitions and examples from our error taxonomy into the prompt is critical to achieve high precision with BooookScore. [^9] In preliminary experiments without definitions and few-shot demonstrations, we qualitatively observe significantly reduced annotation precision.
@@ 4.1 Implementing BooookScore
Assume we have a summary $S$ consisting of sentences $s_{1},s_{2},\dots,s_{n}$. [^10] After iterating over the design in numerous preliminary experiments, we find that our prompt works most reliably at the sentence level, rather than at the full summary level. As such, sentence tokenization is a required preprocessing step for BooookScore. Future work should focus on implementations at the summary level, as it would save many calls to the LLM; here, we need to prompt the model separately for each sentence. We develop a few-shot error-identification prompt $E$ that instructs the LLM to identify any instances of one of the eight specified error types in a given sentence $s_{i}$ of the summary. Concretely, we iterate over each sentence $s_{i}$ in the summary, feeding the prompt $E$, full summary $S$, and target sentence $s_{i}$ at each step. There are two acceptable outputs at each step: either (1) no error is found and the LLM outputs No confusion, or (2) an error(s) is identified and the LLM is asked to generate a corresponding question and associated error type. We include two full summaries with 42 sentence-level annotations in the prompt as demonstrations. [^11] These examples contain a combination of sentences with and without confusion, all the while maintaining a diverse range of error types. The full prompt can be found in J.4. The BooookScore of a single summary $S$ (Figure 2) is then computed as:
@@
We validate BooookScore annotations in the same way that we validate human annotations in Section 3: by hiring human annotators to judge whether they agree with an LLM-generated annotation (here, GPT-4). We observe that the precision of human annotations is 79.7%, while the precision of BooookScore annotations is 78.2% (details in Appendix G). Additionally, we compute BooookScore using human annotations instead of LLM-generated ones for both GPT-4 configurations (i.e., replacing $\text{LLM}(E,S,s_{i})$ in Equation 1 with the human error annotation for $s_{i}$) and observe extremely similar system-level scores. Using human annotations in Equation 1 yields a BooookScore of 82.1 and 89.4131313Recall that human annotators can (1) highlight multiples consecutive sentences as one span and (2) create relations between two spans, while GPT-4 can only highlight single sentences as spans. To adjust for this difference, we treat both consecutive sentences and relations as single sentences when computing BooookScore for humans. for GPT-4 summaries generated via incremental updating and hierarchical merging, respectively, while using LLM annotations yields a BooookScore of 82.4 and 90.8. Figure 3 compares the error distributions from GPT-4 to those of human annotators and shows that GPT-4 is more sensitive to omission errors and less sensitive to duplication or language errors. Taken as a whole, these results confirm that BooookScore is a reliable annotator of coherence for book-length summarization. While we implement BooookScore with GPT-4 for the remainder of this paper, implementing BooookScore with open-source LLM annotators is an exciting future direction.
@@ 5 Systematic evaluation of LLMs
Armed with BooookScore, we now investigate the impact of several critical implementation decisions on summary coherence, including the choice of prompting strategy, base LLM, and chunk size. Overall, Claude 2 produces the most coherent summaries as measured by BooookScore, followed closely by GPT-4 and distantly by ChatGPT and LLaMA 2; however, GPT-4’s summaries are significantly longer and more detailed than the others across both prompting strategies. The rest of this section drills down into finer-grained results.
@@
Table 2 contains results for four instruction-tuned LLMs: GPT-4, ChatGPT, Claude 2, and LLaMA 2.141414Specifically, we use the gpt-4 2023-03-15 and gpt-3.5-turbo-0301 checkpoints on Microsoft Azure. Anthropic unfortunately does not disclose checkpoint information, but our summaries were all obtained via their Claude 2 API in September 2023. We use the LLaMA-2-7B-Instruct checkpoint [^15] https://github.com/togethercomputer/Llama-2-7B-32K-Instruct fine-tuned on long-context summarization (BookSum) and later describe failed experiments with other LLaMA 2 checkpoints. Unless otherwise specified, we set the chunk size to 2048, maximum summary length $G_{n}$ to 900, decoding temperature to 0.5, [^16] Claude 2 is the only exception, as we use its default temperature of 1. and $p=1$ for nucleus sampling. [^17] We use a temperature of 1 for compression, which improves adherence to the max summary length. GPT-4 configurations in this table are not comparable to the ones we analyzed in Section 3 since we had to reduce chunk size and summary length due to LLaMA2 and ChatGPT’s smaller context size. To avoid confounds, we use identical prompts for all models except LLaMA 2, which only functions with a simpler prompt. LLM API costs for our experiments were \$9K (Table 6); more details in Appendix D.
@@
Hierarchical summaries generally have lower BooookScore than incremental summaries, likely because the incremental updating task requires the base LLMs to follow more complex instructions (e.g., deciding what information to include from the current book chunk, what information to discard from the current summary, and whether to restructure the current summary). While hierarchical summarization potentially drops long-range dependencies, its instructions are generally simpler (summarize or merge). ChatGPT is the worst incremental summarizer: it obtains the lowest BooookScore of all of our configurations (67.0), likely because it is highly extractive which negatively impacts coherence (only 68.2% of its generated trigrams were not present in the source).
@@
LLaMA 2 struggles on book-length summarization.
Table 2 shows that LLaMA 2 achieves by far the worst hierarchical BooookScore of any model. This is also the only configuration whose summaries contain significant repetition (% of repeated trigrams), which is a critical coherence error. Furthermore, we could not get the LLaMA-2-7B-Instruct checkpoint to perform incremental updating at all, as it just copied text from the chunks until it reached the summary length limit, at which point it failed to follow the compression instruction. As mentioned before, we also tried both prompting strategies with the larger LLaMA-2-70B-Chat checkpoint, but this model could not even generate hierarchical summaries, often choosing to regurgitate the prompt instead. While future work should evaluate other open-source LLMs, our experiments suggest closed-source models have much more powerful instruction-following abilities for book-length summarization.
@@
Appendix I contains summaries generated from Janika Oz’s A History of Burning, which tells a multi-generational story about an Indian family living in Uganda. We observe that both GPT-4 and ChatGPT tend to generate oft-repetitive and vague sentences within their summaries (e.g., The story highlights the resilience and determination of the characters as they navigate the complexities of life, love, and identity across generations and continents.). Such artifacts are rarely produced by the 88K chunk size version of Claude 2, which instead omits key information present in the beginning or middle of the input (e.g., the entire story of the first generation in the book) in favor of focusing on the end of the book, following the findings of Liu et al. (2023a). All configurations make faithfulness errors: for example, in A History of Burning, the mother of the character Hari is incorrectly identified as Rajni by Claude 2, while GPT-4 does describe Hari’s parentage correctly at one point in the summary but incorrectly at another. We show in Appendix F that automatic quality metrics such as BLANC (Vasilyev et al., 2020) and SUPERT (Gao et al., 2020) are inadequate for book-length summarization.

### v3

@@ Abstract
Summarizing book-length documents (>>100K tokens) that exceed the context window size of large language models (LLMs) requires first breaking the input document into smaller chunks and then prompting an LLM to merge, update, and compress chunk-level summaries. Despite the complexity and importance of this task, it has yet to be meaningfully studied due to the challenges of evaluation: existing book-length summarization datasets (e.g., BookSum) are in the pretraining data of most public LLMs, and existing evaluation methods struggle to capture errors made by modern LLM summarizers. In this paper, we present the first study of the coherence of LLM-based book-length summarizers implemented via two prompting workflows: (1) hierarchically merging chunk-level summaries, and (2) incrementally updating a running summary. We obtain 1193 fine-grained human annotations on GPT-4 generated summaries of 100 recently-published books and identify eight common types of coherence errors made by LLMs. Because human evaluation is expensive and time-consuming, we develop an automatic metric, BooookScore, that measures the proportion of sentences in a summary that do not contain any of the identified error types. BooookScore has high agreement with human annotations and allows us to systematically evaluate the impact of many other critical parameters (e.g., chunk size, base LLM) while saving \$15K USD and 500 hours in human evaluation costs. We find that closed-source LLMs such as GPT-4 and Claude 2 produce summaries with higher BooookScore than those generated by open-source models. While LLaMA 2 falls behind other models, Mixtral achieves performance on par with GPT-3.5-Turbo. Incremental updating yields lower BooookScore but higher level of detail than hierarchical merging, a trade-off sometimes preferred by annotators. We release code and annotations to spur more principled research on book-length summarization.
[Uncaptioned image] github.com/lilakk/BooookScore
@@ 1 Introduction
Despite the promise that LLMs hold for long-context tasks, the research community still lacks a principled and systematic approach to evaluate their capabilities on book-length summarization. Our paper identifies three open challenges with evaluation: (1) data contamination, in which existing benchmarks such as BookSum (Kryscinski et al., 2022) are in the pretraining data of modern LLMs (Chang et al., 2023); (2) an unexplored error distribution, as most prior summarization research centers around short source documents and fails to capture coherence errors that are exacerbated by the “chunk and combine” book-length summarization setting; and (3) a lack of any reliable automatic metric, which requires careful design and validation against human annotations.
@@
To mitigate the impact of data contamination, we design our evaluation framework around the use of newly-published books. We propose a reference-free protocol that leverages human annotation of the coherence of LLM-generated summaries (i.e., their logical connectedness) under different prompting strategies. Our protocol unifies and extends best-practices across disparate works in document understanding and evaluation research, including adoption of fine-grained annotation units, use of QA pairs to denote points of confusion, and a taxonomic breakdown of different coherence errors.
We validate our protocol by collecting 1193 span-level human annotations on GPT-4 generated summaries of a carefully curated set of 100 recently-published books (costing \$3K USD and 100 annotator hours) using two prompting strategies (hierarchical merging and incremental updating, shown in Figure 1). In categorizing these annotations into eight frequent error types, we reveal an error distribution in GPT-4 summaries that differs from that observed in prior studies on short-document summarizers; notably, we identify new error types (causal omissions, salience errors) through our book-length summarization setting (Table 3).
@@
We use BooookScore to evaluate the impact of several critical design decisions on the coherence of generated summaries, including the choice of prompting strategy, base LLM, and chunk size, a study that altogether cost \$10K (USD) in LLM API calls. Our findings include (1) hierarchical merging generally results in more coherent summaries but reduced level of detail compared to incremental updating; (2) GPT-4 and Claude 2 produce the most coherent summaries, while LLaMA 2 is substantially worse and fails to follow instructions; (3) increasing the chunk size does not improve hierarchical merging but does substantially benefit Claude 2 when using incremental updating; and (4) summary-level preference judgments are highly subjective and do not correlate with BooookScore.
@@ 2 Background: summarizing book-length texts with LLMs
It is possible that since hierarchical merging necessitates summarizing portions of the input document without complete context, it may introduce more coherence errors. For example, in the first level, chunks towards the end of the book will be summarized without knowledge of what came before, which can lead to incoherent summaries especially for non-linear or multi-perspective narratives. We thus explore an alternate prompting strategy—incremental updating (Figure 1, right)— that iterates through each chunk in order while continuously updating a global summary with salient information. While this method may be better able to handle inter-chunk dependencies than hierarchical merging, it requires more complicated prompts for (1) summarizing an input chunk, (2) updating the global summary $s_{1,2,\dots,i-1}$ with information from the current chunk $c_{i}$, and (3) compressing the global summary when it exceeds the maximum summary length $G_{n}$. See Appendix A.2 for a full specification of incremental updating.
@@ 3 Evaluating coherence of book summaries
The only existing public dataset for book-length summarization is BookSum (Kryscinski et al., 2022), which contains famous books from the Project Gutenberg public-domain repository along with reference summaries scraped from popular websites such as CliffNotes and GradeSaver. Both the source books and reference summaries are in the pretraining data of existing LLMs: Chang et al. (2023) confirm that many books in the BookSum held-out split (e.g., The Adventures of Huckleberry Finn, The Picture of Dorian Gray) are among the most-memorized books by GPT-4 and GPT-3.5-Turbo, and we were able to auto-complete several reference BookSum summaries by prompting GPT-4 with a short prefix of the summary.
To reduce the confounding impact of summary memorization, we manually collect 100 books [^3] We roughly balance our dataset across the following genres: fiction, non-fiction, sci-fi, fantasy, historical, contemporary, and memoir. We also include both linear and non-linear (multi-perspective and time-shifting) narratives in the dataset, and we purchase electronic copies of each of the 100 books in the dataset. published within the past year to form our dataset (see Table 3 for a full list). Some of these books could still have appeared in the pretraining dataset of recent LLMs such as Claude 2 and LLaMa2, although it is much less likely than in BookSum. However, summaries of these books do not publicly exist: we did not find summaries online for any books in our dataset, which significantly lowers the possibility of LLM memorization. [^4] However, we did find book reviews, which intentionally do not reveal major plot points or other spoilers. The average length of the books in our dataset is 190K tokens, compared to 112K tokens in BookSum. Due to copyright laws, we cannot publicly release this dataset; even if we could, we would still recommend that researchers collect their own datasets of newly-published books to minimize contamination with LLMs of the future.
@@
Since we lack gold summaries, we design our evaluation framework to be reference-free, which aids in scalability. To do this, our evaluation framework synthesizes best-practices of prior document understanding and summarization evaluation research. Our evaluation employs: (1) fine-grained evaluation units as recommended by LongEval (Krishna et al., 2023); (2) information-seeking questions to represent naturally-occurring points of confusion (Ko et al., 2020; Wu et al., 2023; Meng et al., 2023; Newman et al., 2023); and (3) focus on summary coherence, which evaluates the logical structure and readability of the summary itself (Goyal et al., 2022a). We do not directly evaluate the faithfulness of the summaries (i.e., how factually accurate they are at conveying information from the source text), as the length of the source texts poses considerable issues for any existing faithfulness evaluation. We qualitatively discuss faithfulness in Section 5 and leave further investigation for future work.
@@
We implement our framework through a source- and reference-free annotation protocol where (1) annotators read through an LLM-generated summary, (2) highlight all confusing spans, and (3) ask question(s) for each marked span that highlight their confusion. [^5] We also enabled forming relations between two spans in case multiple spans contributed to the same issue. See Table 3 (third column) for examples of spans and questions produced by our annotators. We hired four annotators with extensive English proofreading experience on Upwork [^6] http://upwork.com, each of whom annotated 25 disjoint summaries. Each summary takes roughly 30 minutes to fully annotate with spans and questions, and we paid \$15 USD per summary for a total of \$3K to evaluate both prompting strategies. To generate the summaries, we set the base LLM to GPT-4 with a chunk size of 4096 and a maximum summary length $G_{n}=1200$; other hyperparameters are detailed in Section 5. In total, the annotators mark 840 (incremental updating) and 353 (hierarchical merging) coherence errors for GPT-4-generated summaries; see Table 3 (right) for the split across error types.
@@
Typical measures of agreement are difficult to obtain in our setup, as measuring recall would require ground truth annotations with all possible coherence errors in the summaries; additionally, Goyal et al. (2022a) and  Dou et al. (2022) observed low recall among annotators when evaluating machine-generated text at a fine-grained level. This motivates us to instead measure the precision of a given error annotation (i.e., after reading the corresponding question, do you agree that the span is confusing?), as it is simpler and cheaper while still being an informative metric. Given a span from a summary marked as containing an error, along with questions highlighting the confusion, we ask annotators (1) whether they think the span is confusing; and (2) whether the corresponding questions highlight the central confusion. We use the same four annotators hired before for this task, but make them validate human and (and later GPT-4) annotations for 25 books that they did not annotate in the first task. Overall, we validated 1,659 annotations for a total cost of \$418.90 (USD), [^7] This cost includes validation of both human and BooookScore annotations. and we discover that 79.7% of annotated spans are validated as legitimate through this task. More details on our validation can be found in Appendix J.
@@
Table 1: Definition of all coherence error types, an example annotation for each, and their prevalence (%) in generated summaries, which is calculated as the number of error occurrences in all summaries normalized by the total number of sentences in all summaries.
@@ 4 BooookScore: an automatic evaluation metric
Since human evaluation of summary coherence is not scalable due to the high financial and time cost, we develop an automatic metric — BooookScore— that prompts an LLM to identify instances of the eight error types we identified in Section 3. We validate BooookScore via a human evaluation of its precision (following the annotation task discussed in the previous section) and show that its precision matches that of human annotators (78.2% vs. 79.7%). We then use BooookScore to evaluate many other book-length summarization configurations, saving \$15K USD in evaluation costs and 500 hours in annotator time. We emphasize that incorporating definitions and examples from our error taxonomy into the prompt is critical to achieve high precision with BooookScore. [^8] In preliminary experiments without definitions and few-shot demonstrations, we qualitatively observe significantly reduced annotation precision.
@@ 4.1 Implementing BooookScore
Assume we have a summary $S$ consisting of sentences $s_{1},s_{2},\dots,s_{n}$. [^9] After iterating over the design in numerous preliminary experiments, we find that our prompt works most reliably at the sentence level, rather than at the full summary level. As such, sentence tokenization is a required preprocessing step for BooookScore. Future work should focus on implementations at the summary level, as it would save many calls to the LLM; here, we need to prompt the model separately for each sentence. We develop a few-shot error-identification prompt $E$ that instructs the LLM to identify any instances of one of the eight specified error types in a given sentence $s_{i}$ of the summary. Concretely, we iterate over each sentence $s_{i}$ in the summary, feeding the prompt $E$, full summary $S$, and target sentence $s_{i}$ at each step. There are two acceptable outputs at each step: either (1) no error is found and the LLM outputs No confusion, or (2) an error(s) is identified and the LLM is asked to generate a corresponding question and associated error type. We include two full summaries with 42 sentence-level annotations in the prompt as demonstrations. [^10] These examples contain a combination of sentences with and without confusion, all the while maintaining a diverse range of error types. The full prompt can be found in M.4. The BooookScore of a single summary $S$ (Figure 2) is then computed as:
@@
We validate BooookScore annotations in the same way that we validate human annotations in Section 3: by hiring human annotators to judge whether they agree with an LLM-generated annotation (here, GPT-4). We observe that the precision of human annotations is 79.7%, while the precision of BooookScore annotations is 78.2% (details in Appendix J). Additionally, we compute BooookScore using human annotations instead of LLM-generated ones for both GPT-4 configurations (i.e., replacing $\text{LLM}(E,S,s_{i})$ in Equation 1 with the human error annotation for $s_{i}$) and observe extremely similar system-level scores. Using human annotations in Equation 1 yields a BooookScore of 82.1 and 89.4121212Recall that human annotators can (1) highlight multiples consecutive sentences as one span and (2) create relations between two spans, while GPT-4 can only highlight single sentences as spans. To adjust for this difference, we treat both consecutive sentences and relations as single sentences when computing BooookScore for humans. for GPT-4 summaries generated via incremental updating and hierarchical merging, respectively, while using LLM annotations yields a BooookScore of 82.4 and 90.8. Figure 4 compares the error distributions from GPT-4 to those of human annotators and shows that GPT-4 is more sensitive to omission errors and less sensitive to duplication or language errors. Taken as a whole, these results confirm that BooookScore is a reliable annotator of coherence for book-length summarization. While we implement BooookScore with GPT-4 for the remainder of this paper, implementing BooookScore with open-source LLM annotators is an exciting future direction.
@@ 5 Systematic evaluation of LLMs
Armed with BooookScore, we now investigate the impact of several critical implementation decisions on summary coherence, including the choice of prompting strategy, base LLM, and chunk size. Overall, Claude 2 produces the most coherent summaries as measured by BooookScore, followed closely by GPT-4 and distantly by GPT-3.5-Turbo, Mixtral-8x7B, and LLaMA2-7B-Inst; however, GPT-4’s summaries are significantly longer and more detailed than the others across both prompting strategies. The rest of this section drills down into finer-grained results.
@@
Table 2 contains results for five instruction-tuned LLMs: GPT-4, GPT-3.5-Turbo, Claude 2, Mixtral-8x7B, and LLaMA2-7B-Instruct. [^13] GPT-4 configurations in this table are not comparable to the ones we analyzed in Section 3 since we had to reduce chunk size and summary length due to LLaMA2-7B-Inst and GPT-3.5-Turbo’s smaller context size. Unless otherwise specified, we set the chunk size to 2048, maximum summary length $G_{n}$ to 900, decoding temperature to 0.5, [^14] Claude 2 is the only exception, as we use its default temperature of 1. and $p=1$ for ancestral sampling. [^15] We use a temperature of 1 for compression, which improves adherence to the max summary length. To avoid confounds, we use identical prompts for all models except LLaMA2-7B-Inst, which only functions with a simpler prompt. LLM API costs for our experiments were \$10K USD (Table 8); more experimental details are in Appendix D.
@@
Hierarchical summaries generally have higher BooookScore than incremental summaries, likely because the incremental updating task requires the base LLMs to follow more complex instructions (e.g., deciding what to include from the current book chunk, what to discard from the summary, whether to restructure the summary, etc.). While hierarchical summarization potentially drops long-range dependencies, its instructions are generally simpler (summarize or merge).
@@
LLaMA 2 struggles on book-length summarization while Mixtral shows promising performance.
Table 2 shows that LLaMA-2-7B-Instruct achieves by far the worst hierarchical BooookScore of any model. Its summaries also contain significant repetition (% of repeated trigrams), which is a critical coherence error. Furthermore, we could not get the LLaMA-2-7B-Instruct checkpoint to perform incremental updating at all, as it just copied text from the chunks until it reached the summary length limit, at which point it failed to follow the compression instruction. On the positive side, Mixtral-8x7B, another open-source LLM, outperforms LLaMA-2-7B-Instruct by a substantial margin, though it still trails behind most of the closed-source models. Nonetheless, it is encouraging to note that with performances closely matching that of GPT-3.5-Turbo on both summarization approaches, Mixtral-8x7B signals the narrowing gap between open-source and closed-source models.
@@
Appendix L contains summaries generated from Janika Oz’s A History of Burning, which tells a multi-generational story about an Indian family living in Uganda. We observe that both GPT-4 and GPT-3.5-Turbo tend to generate oft-repetitive and vague sentences within their summaries (e.g., The story highlights the resilience and determination of the characters as they navigate the complexities of life, love, and identity across generations and continents.). Such artifacts are rarely produced by the 88K chunk size version of Claude 2, which instead omits key information present in the beginning or middle of the input (e.g., the entire story of the first generation in the book) in favor of focusing on the end of the book, following the findings of Liu et al. (2023a). All configurations make faithfulness errors: for example, in A History of Burning, the mother of the character Hari is incorrectly identified as Rajni by Claude 2, while GPT-4 does describe Hari’s parentage correctly at one point in the summary but incorrectly at another. We show in Appendix I that automatic quality metrics such as BLANC (Vasilyev et al., 2020) and SUPERT (Gao et al., 2020) are inadequate for book-length summarization.